Problem (2), 10 points Let n be an integer. Prove that if 3 does not divide n, then 3(2n2 5)
Answer #1

Since n is not divisible by 3, n can be represented as 3m + 1 or 3m + 2.

Case 1: n = 3m + 1

2n2 - 5 = 2(3m+1)2- 5 = 2(9m2 + 6m + 1) - 5 = 18m2 + 12m - 3 = 3(6m2 + 4m - 1)

This proves that 3 | 2n2 - 5

Case 2: n = 3m + 1

2n2 - 5 = 2(3m+2)2- 5 = 2(9m2 + 12m + 4) - 5 = 18m2 + 24m + 3 = 3(6m2 + 4m + 1)

This proves that 3 | 2n2 - 5